# Co-Reinforcement Learning for Unified Multimodal Understanding and Generation

### v2

@@ 3.3.2 Unified Reinforcement Learning for Synergistic Multimodal Modeling
As illustrated in Figure 2, the policy ULM first undergoes unified reinforcement learning with diverse rewards across understanding and generation tasks. This unified process aims to jointly enhance its dual capabilities and establish a solid foundation for subsequent task-specific refinement.
@@
where $\lambda$ is a coefficient that balances the two types of rewards. During training, given an input prompt and an image-question pair, the policy model $\pi_{\bm{\theta}_{\mathrm{old}}}$ first generates $G$ candidate responses, $o=\{(\mathcal{I}_{1},\mathcal{T}_{1}),(\mathcal{I}_{2},\mathcal{T}_{2}),\dots,(\mathcal{I}_{G},\mathcal{T}_{G})\}$, each comprising a synthesized image $\mathcal{I}$ and a CoT-format solution $\mathcal{T}$. Concurrently, the joint reward function $\mathcal{R}_{\text{Uni-S1}}$ evaluates each candidate pair, yielding the reward set $r=\{r_{1},r_{2},\dots,r_{G}\}$. These rewards are subsequently normalized according to Eq. 1 to compute the corresponding group-relative advantages $A=\{A_{1},A_{2},\dots,A_{G}\}$. The new policy model $\pi_{\bm{\theta}}$ is then updated by maximizing the following GRPO-based objective:
@@
Notably, based on empirical findings from recent work [87], we omit the KL-divergence constraint during this stage to improve both optimization efficiency and generalization capability.
Training Data.  To support unified RL for synergistic multimodal modeling, we curate a comprehensive dataset comprising 22K samples [^2] https://huggingface.co/datasets/mm-vl/x2x_rft_22k, which follows the data structure established in Sec. 3.2. Each sample includes a real image, a prompt for visual generation, and a CoT-format QA pair for multimodal understanding. This balanced data composition facilitates joint optimization of dual capabilities within a unified framework, while preserving the granularity of task-specific supervision.
@@ 3.3.3 Refined Reinforcement Learning for Task-specific Enhancement
After completing unified RL, as shown in Figure 2, we apply a targeted learning strategy to further enhance the task-specific performance of the policy model. This second-stage optimization leverages task-specific rewards and tailored datasets for individual tasks.
Reward Function and Training Objective.  For text-to-image generation, the reward is defined as $\mathcal{R}_{\text{T2I-S2}}=\mathcal{R}_{\text{cycle}}+\mathcal{R}_{\text{TIM}}$. For multimodal understanding, we define two distinct reward formulations: (1) $\mathcal{R}_{\text{MCQ-S2}}=\mathcal{R}_{\text{MCQ-Acc}}+\mathcal{R}_{\text{Format}}$ for multiple-choice questions, and (2) $\mathcal{R}_{\text{OE-S2}}=\mathcal{R}_{\text{OE-Acc}}+\mathcal{R}_{\text{Format}}$ for open-ended questions. The training objective in this stage adheres to the standard GRPO formulation in Eq. 2, with the appropriate task-specific reward ($\mathcal{R}_{\text{T2I-S2}}$, $\mathcal{R}_{\text{MCQ-S2}}$, or $\mathcal{R}_{\text{OE-S2}}$) replacing $A_{i}$ depending on the task. To ensure stable optimization, we reintroduce the KL-divergence constraint at this stage to limit policy deviation from the reference distribution.
Training Data.  For text-to-image generation, we continue training on the curated dataset introduced in Sec. 3.2. For multimodal understanding, we utilize two specialized datasets: mcot_r1_mcq [^3] https://huggingface.co/datasets/mm-vl/mcot_r1_mcq_66k for multiple-choice questions and mcot_r1_vqa [^4] https://huggingface.co/datasets/mm-vl/mcot_r1_vqa_66k for open-ended questions. These task-specific datasets enable the model to develop more refined and robust capabilities within each task domain.
@@ 4 Experiment
PixArt-$\alpha$ [4] 0.6B 5122 Diff 0.50 0.44 0.08 0.07 0.48 0.47 71.11
SDv1.5 [56] 0.9B 5122 Diff 0.38 0.35 0.04 0.06 0.43 0.32 63.18
SDv2.1 [56] 0.9B 5122 Diff 0.51 0.44 0.07 0.17 0.50 0.32 68.09
SD3-Medium [14] 2B 5122 Diff 0.94 0.72 0.33 0.60 0.74 0.42 84.08
SDXL [51] 2.6B 10242 Diff 0.74 0.39 0.15 0.23 0.55 0.43 74.65
DALL·E 3 [3] - 10242 Diff 0.87 0.47 0.43 0.45 0.67 - 83.50
LlamaGen [62] 0.8B 2562 F-AR 0.34 0.21 0.07 0.04 0.32 - 65.16
@@
LWM [36] 7B 5122 F-AR 0.41 0.46 0.09 0.15 0.47 - -
Orthus [26] 7B 5122 AR-Diff - - - - 0.58 0.27 -
Janus-Pro [7] 7B 3842 F-AR 0.89 0.59 0.79 0.88 0.80 0.35 84.19
ILLUME+ [21] 3B 3842 AR-Diff 0.88 0.62 0.42 0.53 0.72 - -
D-DiT [34] 2B 5122 Diff 0.80 0.54 0.32 0.50 0.65 - -
Harmon [80] 1.5B 5122 F-AR 0.86 0.66 0.74 0.48 0.76 0.41 -
show-o [83] 1.3B 5122 AR-Diff 0.80 0.66 0.31 0.50 0.68 0.35 67.48
@@
Janus [78] 1.3B 3842 F-AR 0.68 0.30 0.46 0.42 0.61 0.23 79.68
Janus-Pro [7] 1.5B 3842 F-AR 0.82 0.51 0.65 0.56 0.73 0.26 82.63
@@ 4.1 Experimental Setups
Implementation Details.  We develop ULM-R1 using Janus-Pro-1B [7] as the baseline ULM for unified multimodal understanding and generation. To ensure reproducibility and scalability, our RL training is built upon the trl [68] framework. In the unified RL stage, we employ the AdamW optimizer with an initial learning rate of 4e-6 and a batch size of 16. We sample 8 responses for both understanding and generation tasks, and set the reward balancing factor in Eq. 5 to 0.8. In the refined RL stage, we sample 16 responses for both multimodal understanding and text-to-image generation tasks. Additionally, we reduce the learning rate to 1e-6 to facilitate fine-grained optimization. All training is conducted on 8 NVIDIA H20 (96G) GPUs. During inference, greedy decoding is used for text generation in multimodal understanding tasks. For text-to-image generation, we employ classifier-free guidance (CFG) [19] with a guidance weight set to 5. More details on the training data and settings are provided in App. A.
@@
SmolVLM [45] SmolLM2-1.7B 38.8 41.7 9.1 33.8 85.5 28.0 43.6 12.6 12.8
SAIL-VL [10] Qwen2.5-1.5B 44.1 56.5 14.6 44.2 88.1 30.4 62.8 17.4 17.3
Ovis2 [42] Qwen2.5-1.5B 45.6 56.7 9.9 58.3 87.8 34.7 64.1 29.4 17.7
InternVL3 [103] Qwen2.5-1.5B 48.7 61.1 22.9 67.0 90.1 34.7 57.6 24.5 20.2
Qwen2.5-VL [2] Qwen2.5-3B 51.2 56.3 22.9 60.0 85.9 40.3 61.2 31.2 21.9
LMM-R1 [49] Qwen2.5-3B - 58.0 - - - - 63.2 41.6 26.4
@@
ILLUME+ [21] Qwen2.5-3B 44.3 - - 40.3 87.6 - - - -
Harmon [80] Qwen2.5-1.5B 38.9 - - - 87.6 - - - -
VILA-U [81] LLaMA-2-7B - - - 33.5 85.8 - - - -
Orthus [26] Chameleon-7B 28.2 - - - 79.6 - - - -
UniToken [24] Chameleon-7B 32.8 46.1 - - - - 38.5 -
SGen-VL [28] InternLM2-1.8B 34.2 - - 34.5 85.3 - 42.7 -
Show-o [83] Phi-1.3B 26.7 - - - 80.0 - - - -
HermesFlow [85] Phi-1.3B 28.3 - - - 81.4 - - - -
Janus-Pro [7] DeepSeek-LLM-7B 41.0 46.5 9.7 50.0 87.4 28.0 42.5 15.9 14.7
Janus [78] DeepSeek-LLM-1.3B 30.5 37.6 3.4 34.3 87.0 23.9 33.7 14.9 13.4
Janus-Pro [7] DeepSeek-LLM-1.5B 36.3 43.1 5.9 39.8 86.2 23.9 37.3 13.5 13.4
@@ 4.2 Quantitative Results
Text-to-Image Generation.  Table 1 presents a comprehensive comparison between ULM-R1 and state-of-the-art models across three visual generation benchmarks. Among unified models, our model ranks second on both GenEval and WISE benchmarks. Notably, it achieves balanced performance across diverse task categories within GenEval, with the best score of 0.71 in object counting. When compared with specialized generation-only models, ULM-R1 surpasses the top performer SD3-Medium [14] by a slight margin (0.77 vs. 0.74 on GenEval). Moreover, ULM-R1 shows consistent improvements over its base model across all benchmarks. These results collectively demonstrate the effectiveness and advantage of our CoRL in enhancing visual generation quality.
Multimodal Understanding.  Results are shown in Table 2. We continue to apply the Gaussian distribution-based merging strategy [61] to combine the two task-specific policy models for mixed QA format evaluation. Overall, ULM-R1 markedly outperforms existing unified models across most benchmarks, and substantially narrows the performance gap with leading understanding-only MLLMs of comparable model scale. More specifically, our model achieves state-of-the-art performance among unified models on MMStar (47.6), WeMath (21.1), LogicVista (34.5), and on several mixed-format math benchmarks, including MathVerse (25.4) and MathVision (22.0). Particularly, ULM-R1 demonstrates considerable improvements over its base model in mathematical and logical reasoning tasks, achieving gains of 15.2 on WeMath and 10.6 on LogicVista. These results not only demonstrate the effectiveness of CoRL in enhancing ULMs’ understanding capabilities, but also establish that reinforcement learning provides a data-efficient pathway for achieving both robust generalization and sophisticated reasoning capabilities, without the need for large-scale supervised data.
@@ 4.3 Qualitative Results
In this section, we first present a qualitative comparison between ULM-R1 and Janus-Pro on visual generation, as illustrated in Figure 3. The results clearly showcase that ULM-R1 achieves superior text-to-image alignment and object grounding across diverse prompts, with especially notable improvements in spatial object arrangement and compositional consistency. Furthermore, as shown in Figure 4, we visualize several representative examples of multimodal understanding. Compared to Janus-Pro, ULM-R1 exhibits significantly enhanced understanding capabilities, particularly in mathematical reasoning. These comprehensive qualitative results demonstrate the effectiveness of CoRL in simultaneously improving both visual generation and multimodal understanding in ULMs.
4.4 Ablation Studies
In this section, we primarily validate the effectiveness of our training strategy and designed rewards.
@@ 4.4 Ablation Studies
# $\mathcal{R}_{\mathrm{cycle}}$ $\mathcal{R}_{\mathrm{TIM}}$ GenEval DPG Avg. $\uparrow$
0 73.0 82.6 77.8
1 76.2 83.5 79.9 (+2.1)
2 74.1 83.0 78.6 (+0.8)
3 77.3 83.9 80.6 (+2.8)
Rewards in Text-to-Image Generation.  To assess the effectiveness of our proposed rewards for text-to-image generation, we conduct ablation experiments as detailed in Table 4. The results demonstrate that incorporating either reward individually improves performance over the baseline: $\mathcal{R}_{\mathrm{cycle}}$ yields an increase of 2.1 in average score, while $\mathcal{R}_{\mathrm{TIM}}$ results in an increase of 0.8. Notably, combining both rewards leads to the best overall performance, achieving an average score of 80.6. These findings suggest a modest but complementary effect between $\mathcal{R}_{\mathrm{cycle}}$ and $\mathcal{R}_{\mathrm{TIM}}$, enhancing their joint benefit in enhancing visual generation quality.

### v3

@@ 3.3.2 Unified Reinforcement Learning for Synergistic Multimodal Modeling
As illustrated in Figure˜2, the policy ULM first undergoes unified reinforcement learning with diverse rewards across understanding and generation tasks. This unified process aims to jointly enhance its dual capabilities and establish a solid foundation for subsequent task-specific refinement.
@@
where $\lambda$ is a coefficient that balances the two types of rewards. During training, given an input prompt and an image-question pair, the policy model $\pi_{\bm{\theta}_{\mathrm{old}}}$ first generates $G$ candidate responses, $o=\{(\mathcal{I}_{1},\mathcal{T}_{1}),(\mathcal{I}_{2},\mathcal{T}_{2}),\dots,(\mathcal{I}_{G},\mathcal{T}_{G})\}$, each comprising a synthesized image $\mathcal{I}$ and a CoT-format solution $\mathcal{T}$. Concurrently, the joint reward function $\mathcal{R}_{\text{Uni-S1}}$ evaluates each candidate pair, yielding the reward set $r=\{r_{1},r_{2},\dots,r_{G}\}$. These rewards are subsequently normalized according to Eq.˜1 to compute the corresponding group-relative advantages $A=\{A_{1},A_{2},\dots,A_{G}\}$. The new policy model $\pi_{\bm{\theta}}$ is then updated by maximizing the following GRPO-based objective:
@@
Notably, based on empirical findings from recent work [yu2025dapo], we omit the KL-divergence constraint during this stage to improve both optimization efficiency and generalization capability.
Training Data.  To support unified RL for synergistic multimodal modeling, we curate a comprehensive dataset comprising 22K samples [^2] https://huggingface.co/datasets/mm-vl/x2x_rft_22k, which follows the data structure established in Sec.˜3.2. Each sample includes a real image, a prompt for visual generation, and a CoT-format QA pair for multimodal understanding. This balanced data composition facilitates joint optimization of dual capabilities within a unified framework, while preserving the granularity of task-specific supervision.
@@ 3.3.3 Refined Reinforcement Learning for Task-specific Enhancement
After completing unified RL, as shown in Figure˜2, we apply a targeted learning strategy to further enhance the task-specific performance of the policy model. This second-stage optimization leverages task-specific rewards and tailored datasets for individual tasks.
Reward Function and Training Objective.  For text-to-image generation, the reward is defined as $\mathcal{R}_{\text{T2I-S2}}=\mathcal{R}_{\text{cycle}}+\mathcal{R}_{\text{TIM}}$. For multimodal understanding, we define two distinct reward formulations: (1) $\mathcal{R}_{\text{MCQ-S2}}=\mathcal{R}_{\text{MCQ-Acc}}+\mathcal{R}_{\text{Format}}$ for multiple-choice questions, and (2) $\mathcal{R}_{\text{OE-S2}}=\mathcal{R}_{\text{OE-Acc}}+\mathcal{R}_{\text{Format}}$ for open-ended questions. The training objective in this stage adheres to the standard GRPO formulation in Eq.˜2, with the appropriate task-specific reward ($\mathcal{R}_{\text{T2I-S2}}$, $\mathcal{R}_{\text{MCQ-S2}}$, or $\mathcal{R}_{\text{OE-S2}}$) replacing $A_{i}$ depending on the task. To ensure stable optimization, we reintroduce the KL-divergence constraint at this stage to limit policy deviation from the reference distribution.
Training Data.  For text-to-image generation, we continue training on the curated dataset introduced in Sec.˜3.2. For multimodal understanding, we utilize two specialized datasets: mcot_r1_mcq [^3] https://huggingface.co/datasets/mm-vl/mcot_r1_mcq_66k for multiple-choice questions and mcot_r1_vqa [^4] https://huggingface.co/datasets/mm-vl/mcot_r1_vqa_66k for open-ended questions. These task-specific datasets enable the model to develop more refined and robust capabilities within each task domain.
@@ 4 Experiment
PixArt-$\alpha$ [chen2023pixart] 0.6B 5122 Diff 0.50 0.44 0.08 0.07 0.48 0.47 71.11
SDv1.5 [rombach2022high] 0.9B 5122 Diff 0.38 0.35 0.04 0.06 0.43 0.32 63.18
SDv2.1 [rombach2022high] 0.9B 5122 Diff 0.51 0.44 0.07 0.17 0.50 0.32 68.09
SD3-Medium [esser2024scaling] 2B 5122 Diff 0.94 0.72 0.33 0.60 0.74 0.42 84.08
SDXL [podell2023sdxl] 2.6B 10242 Diff 0.74 0.39 0.15 0.23 0.55 0.43 74.65
DALL·E 3 [dalle3] - 10242 Diff 0.87 0.47 0.43 0.45 0.67 - 83.50
LlamaGen [sun2024autoregressive] 0.8B 2562 F-AR 0.34 0.21 0.07 0.04 0.32 - 65.16
@@
LWM [liu2025world] 7B 5122 F-AR 0.41 0.46 0.09 0.15 0.47 - -
Orthus [kou2024orthus] 7B 5122 AR-Diff - - - - 0.58 0.27 -
Janus-Pro [chen2025januspro] 7B 3842 F-AR 0.89 0.59 0.79 0.88 0.80 0.35 84.19
ILLUME+ [huang2025illume+] 3B 3842 AR-Diff 0.88 0.62 0.42 0.53 0.72 - -
D-DiT [li2024dual] 2B 5122 Diff 0.80 0.54 0.32 0.50 0.65 - -
Harmon [wu2025harmonizing] 1.5B 5122 F-AR 0.86 0.66 0.74 0.48 0.76 0.41 -
show-o [xie2025show] 1.3B 5122 AR-Diff 0.80 0.66 0.31 0.50 0.68 0.35 67.48
@@
Janus [wu2024janus] 1.3B 3842 F-AR 0.68 0.30 0.46 0.42 0.61 0.23 79.68
Janus-Pro [chen2025januspro] 1.5B 3842 F-AR 0.82 0.51 0.65 0.56 0.73 0.26 82.63
@@ 4.1 Experimental Setups
Implementation Details.  We develop ULM-R1 using Janus-Pro-1B [chen2025januspro] as the baseline ULM for unified multimodal understanding and generation. To ensure reproducibility and scalability, our RL training is built upon the trl [vonwerra2022trl] framework. In the unified RL stage, we employ the AdamW optimizer with an initial learning rate of 4e-6 and a batch size of 16. We sample 8 responses for both understanding and generation tasks, and set the reward balancing factor in Eq.˜5 to 0.8. In the refined RL stage, we sample 16 responses for both multimodal understanding and text-to-image generation tasks. Additionally, we reduce the learning rate to 1e-6 to facilitate fine-grained optimization. All training is conducted on 8 NVIDIA H20 (96G) GPUs. During inference, greedy decoding is used for text generation in multimodal understanding tasks. For text-to-image generation, we employ classifier-free guidance (CFG) [ho2022classifier] with a guidance weight set to 5. More details on the training data and settings are provided in App.˜A.
@@
SmolVLM [marafioti2025smolvlm] SmolLM2-1.7B 38.8 41.7 9.1 33.8 85.5 28.0 43.6 12.6 12.8
SAIL-VL [dong2025scalable] Qwen2.5-1.5B 44.1 56.5 14.6 44.2 88.1 30.4 62.8 17.4 17.3
Ovis2 [lu2024ovis] Qwen2.5-1.5B 45.6 56.7 9.9 58.3 87.8 34.7 64.1 29.4 17.7
InternVL3 [zhu2025internvl3] Qwen2.5-1.5B 48.7 61.1 22.9 67.0 90.1 34.7 57.6 24.5 20.2
Qwen2.5-VL [bai2025qwen2] Qwen2.5-3B 51.2 56.3 22.9 60.0 85.9 40.3 61.2 31.2 21.9
LMM-R1 [peng2025lmm] Qwen2.5-3B - 58.0 - - - - 63.2 41.6 26.4
@@
ILLUME+ [huang2025illume+] Qwen2.5-3B 44.3 - - 40.3 87.6 - - - -
Harmon [wu2025harmonizing] Qwen2.5-1.5B 38.9 - - - 87.6 - - - -
VILA-U [wu2025vila] LLaMA-2-7B - - - 33.5 85.8 - - - -
Orthus [kou2024orthus] Chameleon-7B 28.2 - - - 79.6 - - - -
UniToken [jiao2025unitoken] Chameleon-7B 32.8 46.1 - - - - 38.5 -
SGen-VL [li2024synergen] InternLM2-1.8B 34.2 - - 34.5 85.3 - 42.7 -
Show-o [xie2025show] Phi-1.3B 26.7 - - - 80.0 - - - -
HermesFlow [yang2025hermesflow] Phi-1.3B 28.3 - - - 81.4 - - - -
Janus-Pro [chen2025januspro] DeepSeek-LLM-7B 41.0 46.5 9.7 50.0 87.4 28.0 42.5 15.9 14.7
Janus [wu2024janus] DeepSeek-LLM-1.3B 30.5 37.6 3.4 34.3 87.0 23.9 33.7 14.9 13.4
Janus-Pro [chen2025januspro] DeepSeek-LLM-1.5B 36.3 43.1 5.9 39.8 86.2 23.9 37.3 13.5 13.4
@@ 4.2 Quantitative Results
Text-to-Image Generation.  Table˜1 presents a comprehensive comparison between ULM-R1 and state-of-the-art models across three visual generation benchmarks. Among unified models, our model ranks second on both GenEval and WISE benchmarks. Notably, it achieves balanced performance across diverse task categories within GenEval, with the best score of 0.71 in object counting. When compared with specialized generation-only models, ULM-R1 surpasses the top performer SD3-Medium [esser2024scaling] by a slight margin (0.77 vs. 0.74 on GenEval). Moreover, ULM-R1 shows consistent improvements over its base model across all benchmarks. These results collectively demonstrate the effectiveness and advantage of our CoRL in enhancing visual generation quality.
Multimodal Understanding.  Results are shown in Table˜2. For mixed QA format evaluation, we continue to apply the Gaussian-distribution-based merging strategy [si2025unveiling] to combine the two task-specific policy models and obtain a final model capable of following both types of instructions. Overall, ULM-R1 markedly outperforms existing unified models across most benchmarks, and substantially narrows the performance gap with leading understanding-only MLLMs of comparable model scale. More specifically, our model achieves state-of-the-art performance among unified models on MMStar (47.6), WeMath (21.1), LogicVista (34.5), and on several mixed-format math benchmarks, including MathVerse (25.4) and MathVision (22.0). Particularly, ULM-R1 demonstrates considerable improvements over its base model in mathematical and logical reasoning tasks, achieving gains of 15.2 on WeMath and 10.6 on LogicVista. These results not only demonstrate the effectiveness of CoRL in enhancing ULMs’ understanding capabilities, but also establish that reinforcement learning provides a data-efficient pathway for achieving both robust generalization and sophisticated reasoning capabilities, without the need for large-scale supervised data.
@@ 4.3 Qualitative Results
In this section, we first present a qualitative comparison between ULM-R1 and Janus-Pro for visual generation, as illustrated in Figure˜3. The results clearly show that ULM-R1 achieves superior text-to-image alignment and object grounding across diverse prompts, with especially notable improvements in spatial arrangement of objects and compositional consistency. Next, as shown in Figure˜4, we visualize several representative examples of multimodal understanding. Compared to Janus-Pro, ULM-R1 exhibits significantly enhanced understanding capabilities, particularly in mathematical reasoning. These comprehensive qualitative results demonstrate the effectiveness of CoRL in simultaneously improving visual generation and multimodal understanding in ULMs. Furthermore, Figures˜3 and 4 respectively showcase exemplary failure cases of ULM-R1 in visual generation and understanding tasks, providing an intuitive grasp of its limitations. For instance, in the understanding example, it misinterprets commonsense and professional knowledge, leading to an incorrect answer.
4.4 Ablation Study and Discussion
In this section, we primarily evaluate the effectiveness of our RL training strategy and the proposed reward functions for text-to-image generation. In addition, we discuss the impact of the hyperparameter $\lambda$ and the scalability of CoRL.
@@ 4.4 Ablation Study and Discussion
Effect of Rewards in Text-to-Image Generation.  To evaluate the effectiveness of our proposed rewards for text-to-image generation, we conduct ablation experiments as detailed in Table˜6. The results demonstrate that incorporating either reward individually improves performance over the baseline: $\mathcal{R}_{\mathrm{cycle}}$ yields an increase of 2.1 in average score, while $\mathcal{R}_{\mathrm{TIM}}$ results in an increase of 0.8. Notably, combining both rewards leads to the best overall performance, achieving an average score of 80.6. These findings suggest a modest but complementary effect between $\mathcal{R}_{\mathrm{cycle}}$ and $\mathcal{R}_{\mathrm{TIM}}$, enhancing their joint benefit in enhancing visual generation quality. In addition, we further compare the CLIP score ($\mathcal{R}_{\mathrm{CLIP}}$) and $\mathcal{R}_{\mathrm{TIM}}$ under our final RL training paradigm. As shown in the table, $\mathcal{R}_{\mathrm{TIM}}$ achieves better overall performance, especially on the DPG benchmark with dense, long-horizon prompts for image generation, highlighting its superior ability to capture fine-grained semantic alignment compared to the CLIP score.
Impact of Visual Consistency Measures in $\mathcal{R}_{\mathrm{cycle}}$.  Table˜6 provides a more detailed analysis of how different visual consistency measures (PSNR, MSE, SSIM, and LPIPS) used in $\mathcal{R}_{\mathrm{cycle}}$ affect the quality of visual generation. PSNR and MSE are pixel-level metrics that quantify low-level differences between the generated images, while SSIM and LPIPS assess higher-level perceptual and structural similarities. As shown in the table, SSIM and LPIPS perform better than the other two metrics, with LPIPS achieving the best performance (83.9) on the DPG benchmark. This can be attributed to the fact that LPIPS measures image similarity in a feature space, making it more robust to minor, semantically irrelevant variations and thus better suited to reward high-level consistency.
@@
Rewards GenEval DPG Avg. $\uparrow$
Baseline 73.0 82.6 77.8
$\mathcal{R}_{\mathrm{CLIP}}$ 74.2 82.4 78.3 (+0.5)
$\mathcal{R}_{\mathrm{TIM}}$ 74.1 83.0 78.6 (+0.8)
$\mathcal{R}_{\mathrm{cycle}}$ 76.2 83.5 79.9 (+2.1)
$\mathcal{R}_{\mathrm{cycle}}$ + $\mathcal{R}_{\mathrm{CLIP}}$ 77.0 83.4 80.2 (+2.4)
$\mathcal{R}_{\mathrm{cycle}}$ + $\mathcal{R}_{\mathrm{TIM}}$ 77.3 83.9 80.6 (+2.8)
Table 5: Comparison among visual consistency measures used in $\mathcal{R}_{\mathrm{cycle}}$.
Measures GenEval DPG
PSNR 76.0 82.4
MSE 77.1 83.2
SSIM 77.5 83.6
LPIPS 77.3 83.9
Table 6: Impact of $\lambda$.
$\lambda$ GenEval MMMU
(Gen.) (Und.)
0.5 77.1 41.0
0.7 77.3 42.3
0.8 77.3 42.3
0.9 77.1 43.5
1.0 76.9 43.0
Impact of hyperparameter $\lambda$.  The factor $\lambda$ in Eq.˜5 balances the reward scales between the two tasks during unified RL. As shown in Table˜6, we conduct experiments using different values of $\lambda$ to assess its impact on both generation and understanding performance. The results show that moderate values of $\lambda$ ($\sim$ 0.8) achieve a balanced trade-off between generation and understanding. Larger values slightly degrade generation performance, indicating that overemphasizing understanding rewards may hinder cross-task optimization.
Scalability of CoRL.  To validate the effectiveness of CoRL on other ULMs, as illustrated in Table˜7, we conduct additional experiments using Janus-1.3B [wu2024janus] and Janus-Pro-7B [chen2025januspro] as the baseline. The results show consistent improvements across both generation and understanding benchmarks, confirming the scalability of CoRL. Notably, Janus-Pro-7B with LoRA tuning achieves smaller gains on the mathematical reasoning benchmark (WeMath) than Janus-1.3B, suggesting that while CoRL scales well across model size, its enhancement of complex reasoning does not scale linearly.
[…]
Table 7: Effectiveness of CoRL on other ULMs. For Janus-Pro-7B, we adopt LoRA tuning [hu2022lora] to enable efficient training and mitigate memory pressure during unified RL.
[…]
Janus-1.3B 0.61 0.23 79.68 30.5 37.6 3.4† 34.3 87.0 23.9
+ CoRL (Full Fine-Tuning) 0.64 0.26 80.92 34.6 41.9 16.4 36.9 88.1 27.0
Janus-Pro-7B 0.80 0.35 84.19 41.0 46.5 9.7 50.0 87.4 28.0
+ CoRL (LoRA Tuning) 0.82 0.41 84.97 44.6 49.5 16.0 52.6 88.0 32.4
5 Limitation
Despite the substantial improvements achieved, several limitations remain that warrant further investigation. First, a notable performance gap still exists between generation and understanding tasks of ULMs. Second, our rewards for multimodal understanding are relatively simple and primary. These limitations highlight the need for more sophisticated RL designs that can further enhance understanding capabilities and narrow the performance gap. We hope our work provides valuable insights for future RL research in ULMs.